# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v2

@@ Abstract
Text-guided diffusion models catalyze a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts. Recent advancements introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pre-trained diffusion models for audio modification. Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy. And the lack of attention control hinders the fine-grained manipulations of music. To counteract these limitations, we introduce the Disentangled Inversion technique, which is designed to disentangle the diffusion process into triple branches, thereby magnifying their individual capabilities for both precise editing and preservation. Furthermore, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention and cross-attention with an additional Harmonic Branch to achieve the desired composition and structural information in the target music. Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing whilst safeguarding structural integrity. To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across 10 distinct editing categories, which facilitates both zero-shot and instruction-based music editing tasks. Our method demonstrates unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques. [^1] Audio samples are available at https://MEDIC-Zero.github.io/. [^2] Benchmark data are available at https://huggingface.co/datasets/liuhuadai/ZoME-Bench.
@@ 1 Introduction
Advancements in text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have markedly enhanced audio generation, leveraging their impressive capability for realistic and varied outputs. These models (Liu et al., 2023b; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes in response to specific textual prompts. Early audio editing strategies rely on training models from scratch or test-time optimization, both of which are hampered by intensive computational demands. Recent advances (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through DDIM (Song et al., 2020) and DDPM (Ho et al., ) inversion techniques, but challenges remain.
@@
Key among these challenges is maintaining editing fidelity - ensuring that the model’s output matches the given instructions - and preserving essential content, a hallmark of effective audio editing that targets only necessary changes. Balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving suboptimal for conditional diffusions (Mokady et al., 2023). Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space, yet at the expense of reduced modification capabilities due to noise reduction.
In this work, we rigorously examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis, suggests that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) further amplifies the accumulated errors. Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has demonstrated its power in achieving high fidelity and essential content preservation. Yet, analogous control mechanisms in music editing are noticeably absent. To bridge this gap, we introduce an innovative Disentangled Inversion Control technique. This technique is comprised of two main components: Harmonic Attention Control and Disentangled Inversion. The former integrates cross-attention and mutual self-attention controls, supplemented by a novel Harmonic Branch, which facilitates precise, fine-grained manipulations for both rigid and non-rigid editing scenarios. The latter component strategically deconstructs the diffusion process into three distinct branches, enabling each to realize its full potential independently.
@@
Acknowledging the lack of standardized benchmarks in the field of audio editing, we introduce ZoME-Bench. The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories. Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing. Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
@@
Our contributions can be summarized as follows. 1) We introduce a novel, tuning-free methodology referred to as Disentangled Inversion Control (DIC), designed to facilitate both consistent manipulations of musical elements and intricate non-rigid editing tasks. 2) An Harmonic Attention Control framework is introduced to unify cross-attention control and self-attention control, which enables both rigid and non-rigid editing. 3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error without optimization by branch disentanglement, aiding in accurately editing the music while preserving the structural information. 4) We introduce a tailored benchmark for music editing, named ZoME-Bench, and Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
@@ 2.1 Text-based Audio editing
The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt. Two primary challenges command intense focus. The first is the fidelity of editing, which pertains to the model’s ability to strictly follow the initial instructions. The second challenge involves essential content preservation, a critical attribute of competent audio editing models that ensures only suitable regional modifications are made during the editing process. Existing methodologies for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction. Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load. The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020) inversion strategies to extract diffusion noise vectors that match the source signal. Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework. In this study, we propose a new inversion technique named Disentangled Inversion Control. This technique aims to achieve accurate editing while preserving structural information.
2.2 Inversion Techniques
@@
The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021). While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023) , its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges. For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during the editing process. Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024).
Recently, optimization-based inversion methods using specific latent variables have gained popularity (Ju et al., 2023; Kawar et al., 2023). These are designed to minimize accumulated errors stemming from the DDIM inversion. Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming. Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
Figure 1: The framework and rationale behind Disentangled Inversion Control. Traditional DDIM inversion introduces incremental errors at each step, leading to only a coarse approximation of the source audio’s distribution. The implementation of classifier-free guidance exacerbates the error accumulation. In contrast, our method intricately separates the diffusion process into a triple-branch system, allowing each branch to fully realize its individual potential.
@@ 3.1 Motivation
Figure 1 and Preliminaries in Appendix 2 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.  Hertz et al. (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$. This sparks a plethora of attention control techniques that direct the image domain’s target denoiser network $\hat{\epsilon}_{\theta}$ to better reflect target prompts. Yet, analogous control mechanisms in music editing are noticeably absent.
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach poised to achieve fine-grained zero-shot music editing. DIC strategically disentangles the diffusion process as a triple-branch, granting each branch to optimize its functionality. At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance. We will first introduce Harmonized Attention Control in Section 3.2 and discuss Disentangled Inversion in Section 3.3.
@@ 3.2 Harmonized Attention Control Framework
The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023a). At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer. The attention mechanism is formalized as:
@@
where $Q$ denotes the query features derived from the audio features, $K$, $V$ represent the key and value features, and $M$ is the attention maps. We explore varying semantic transformations of audio content through bifurcated attention control strategies—cross-attention for rigid changes and mutual self-attention for non-rigid adjustments. Finally, we introduce an intermediate branch to host the desired composition and structural information in the target music.
@@ 3.2.1 Cross-Attention Control
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the modified prompt ${\mathcal{P}}^{*}$. We follow the practice of Prompt-to-Prompt and define CAC as Global Attention Refinement and Local Attention Blend.
@@
For simplificity, we define the process of local editing in Equation 10 as:
@@
$\text{CrossEdit}(M_{t},M^{*}_{t},t)=\begin{cases}\text{Refine}(M^{src},M^{tgt})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
@@ 3.2.3 Harmonic Branch Integretion
The naive combination of cross-attention control and mutual self-attention control sequentially would lead to a sub-optimal outcome in the original dual-branch setup, especially failing the global attention refinement. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music. Our unified framework is detailed in Algorithm 2 and illustrated in Figure 2. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. Following this, cross-attention control is applied on $M_{har}$ and $M_{tgt}$ to refine the semantic information for $M_{tgt}$.

### v3

@@ Abstract
Text-guided diffusion models make a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts. Recent works introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pretrained diffusion models for audio modification. Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy. Moreover, existing editing methods fail to achieve effective complex non-rigid music editing while maintaining essential content preservation and high editing fidelity. To counteract these issues, we introduce the Disentangled Inversion technique to disentangle the diffusion process into triple branches, rectifying the deviated path of the source branch caused by DDIM inversion. In addition, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention control and cross-attention control with an intermediate Harmonic Branch to progressively achieve the desired harmonic and melodic information in the target music. Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing while safeguarding content integrity. To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across ten distinct editing categories. This facilitates both zero-shot and instruction-based music editing tasks. Our method achieves unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques. [^1] The project webpage is available at https://MEDIC-Edit.github.io/.
@@ 1 Introduction
Text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have made great progress in audio generation (Evans et al., 2024a, b), leveraging their impressive capability for realistic and varied outputs. In particular, these models (Liu et al., 2023c; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes for specific textual prompts. Early audio editing strategies rely on training models from scratch (Copet et al., 2023; Agostinelli et al., 2023) or test-time optimization (Paissan et al., 2023; Plitsis et al., 2024), hampered by intensive computational demands. Recent works (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) and Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) inversion techniques, but challenges remain.
@@
Key among these challenges are maintaining fidelity of editing - ensuring the editing aligns with the provided instructions - and essential content preservation, ensuring that the unaltered musical attributes in the target prompts remain unchanged. However, balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving sub-optimal for conditional diffusion models (Mokady et al., 2023). Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space. However, this comes at the expense of reduced modification capabilities due to noise reduction.
In this work, we rigorously examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis indicates that while techniques like DDIM inversion provide an editable foundation for audio synthesis, they lack precision and may compromise the integrity of the original audio. The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently not met during text-conditional editing. Consequently, this leads to distortions during the inversion. And the implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) aims to improve text adherence. However, it inadvertently amplifies the accumulated errors from the inversion process. Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has shown promise in achieving high fidelity and essential content preservation. For instance, MusicMagus (Zhang et al., 2024) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks. Nevertheless, these methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 1. We introduce an innovative Disentangled Inversion Control technique to bridge this gap. This technique has two main components: Harmonic Attention Control and Disentangled Inversion. Cross-attention control (Hertz et al., 2022) and mutual self-attention control (Cao et al., 2023) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively. However, blindly combining these two approaches sequentially for music editing can result in sub-optimal outcomes, particularly in the original dual-branch setup, where it struggles with global attention refinement. To address this challenge, we introduce an intermediate branch called the Harmonic Branch, designed to modify both rigid and non-rigid attributes in music progressively. Furthermore, we disentangle the diffusion process into triple branches, correcting the deviation path caused by CFG in the source branch, which affects the essential content preservation. The other branches remain unchanged to ensure the highest possible edit fidelity.
@@
Due to a lack of standardized benchmarks in audio editing, we introduce ZoME-Bench. The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories across rigid and non-rigid tasks. Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing. Experimental results on ZoME-Bench indicate that MEDIC outperforms baselines, achieving significant improvements in essential content preservation and editing fidelity. Additionally, MEDIC demonstrates state-of-the-art performance in the variable-length music editing settings of the MusicDelta dataset.
@@
Our contributions can be summarized as follows. 1) We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks. 2) An Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing. 3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information. 4) We build a new benchmark for music editing, named ZoME-Bench, which supports both zero-shot and instruction-based music editing.
@@ 2 Related Works
Text-based Audio editing. The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt. Existing methodologies for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction. Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load. The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020; Zhang et al., 2024) inversion strategies to extract diffusion noise vectors that match the source signal. Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework. In this work, we propose a new inversion technique named Disentangled Inversion Control. This technique aims to achieve accurate editing while preserving structural information.
@@
Inversion Techniques. The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021). While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023), its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges. For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during editing. Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024). Optimization-based inversion methods using specific latent variables have recently gained popularity (Ju et al., 2023; Kawar et al., 2023). These are designed to minimize accumulated errors stemming from the DDIM inversion. Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming. Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
Figure 1: Comparisons of our method with two-branch inversion techniques, including DDIM Inversion and MusicMagus. (a) Framework of DDIM Inversion, showing configurations with and without classifier-free guidance. (b) Framework of MusicMagus, which incorporates cross-attention control. (c) Framework of our method, featuring disentangled inversion control. (d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch approach.
@@ 3 Disentangled Inversion Control
3.1 Problem Definition and BenchMark Construction
Despite significant work in text-to-audio generation models, particularly with the emergence of latent diffusion models (LDM), research on zero-shot music editing remains limited. Zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to synthesize the desired music, denoted as ${\bm{x}}^{tgt}_{0}$. This synthesized music should align with the target edited text prompt ${\mathcal{P}}^{*}$, which is directly modified from the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt ${\mathcal{P}}$. We compress source audio signals into latent ${\bm{z}}^{src}_{0}$ for inversion.
To systematically validate our proposed method as a plug-and-play strategy for editing models, compare our method with existing zero-shot music editing methods, and compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset named ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps (Doh et al., 2023), spanning ten different editing types that include both rigid and non-rigid modifications. Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
Additionally, we include annotations relevant to attention control, such as blended words, to facilitate detailed evaluations. Further details about our benchmark can be found in Appendix B.
@@ 3.2 Motivation
Figure 1 and Preliminaries in Appendix A reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision. This potentially compromising the essential content preservation. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.  Hertz et al. (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$. This leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts. Yet, analogous control mechanisms for non-rigid music editing are noticeably limited.
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach to achieve both rigid and non-rigid music editing. DIC strategically disentangles the diffusion process as triple branches, allowing each branch to optimize its functionality. At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance. We will first introduce Harmonized Attention Control in Section 3.3 and discuss Disentangled Inversion in Section 3.4.
@@ 3.3 Harmonized Attention Control Framework
The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023b). At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer. The attention mechanism is formalized as:
@@
where $Q$ denotes the query features derived from the audio features, $K$ and $V$ represent the key and value features, and $M$ is the attention map. We explore varying semantic transformations of audio content through harmonized attention control strategies — cross-attention control for rigid changes and mutual self-attention control for non-rigid adjustments. Finally, we introduce an intermediate branch to host the desired harmonic and melodic information in the target music. The framework of harmonized attention control is depicted in Figure 2.
@@ 3.3.1 Cross-Attention Control
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the target prompt ${\mathcal{P}}^{*}$. We follow the practice of Prompt-to-Prompt (Hertz et al., 2022) and define CAC as Global Attention Refinement and Local Attention Blend.
@@
For simplicity, we define the process of local editing in Equation 5 as:
@@
$\text{CrossEdit}(M^{src},M^{tgt},t)=\begin{cases}\text{Refine}(M^{src}_{t},M^{tgt}_{t})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}_{t}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
@@ 3.3.3 Harmonic Branch Integretion
The naive combination of cross-attention control and mutual self-attention control sequentially would lead to sub-optimal results in the original dual-branch setup, especially failing the global attention refinement. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music.
Our unified framework is detailed in Algorithm 2. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. This lays the formal structure of the target music. Following this, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$. As illustrated in Figure 2, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “violin"), while preserving the rigid content semantics (e.g., “with noise"). The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “with noise").